Paul Coddington

Northeast Parallel Architectures Center 111 College Place Syracuse University Syracuse, NY 13244, U.S.A

Phone: (315) 443-4883 FAX: (315) 443-1973 Email: paulc@npac.syr.edu

July 11, 1994

Selection Committee University of Reading

To whom it may concern,

I was in the same group as James Allwright at Southampton University, and knew something of his work although we did not collaborate on any projects. James has worked with me on a couple of projects at the Northeast Parallel Architectures Center at Syracuse University.

James is very bright and strong academically. I worked with him on a research project on parallel graph coloring algorithms and he did very good work. I cannot comment on his teaching abilities since I have only been involved in his research work, however he does not have particularly good communication skills.

Yours sincerely,

Paul Coddington